Water going over a fall splashes on rocks 20 m below. If 20 kg of water flows over the fall per second, how much force do the rocks exert on the falling water?
Answer #1


using work energy theorem


Work done by the force = change in kinetic energy

W = 0.5*m*v^2


But Work done is W = F*S

F*S = 0.5*m*v^2

but v is the speed of water at the rock = sqrt(2*g*h) = sqrt(2*9.81*20) = 19.8 m/s

then


F*S = 0.5*m*v^2

F*20 = 0.5*20*19.8^2


F = 196 N